What is X-ray density and what is pyknometric density?
If the former is 2.178 and latter is 2.165, then what is the fraction of sites unoccupied in the lattice? (Q: from Solid State)

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
pyknometric density:-Pycnometric density is the closest approximation to true density calculated from the molecular weight and crystalline lattice of a substance.
First density = 2.165 * 10 ^3
Difference of density = 2178-2165 = 13 kg/m 3
Therefore,
fraction of unoccupied site = 13/2178 = 5.96 * 10 -3
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

pyknometric density:-Pycnometric density is the closest approximation to true density calculated from the molecular weight and crystalline lattice of a substance.
First density = 2.165 * 10 ^3
Difference of density = 2178-2165 = 13 kg/m 3
Therefore, fraction of unoccupied site = 13/2178 = 5.96 * 10 -3

Thanks and Regards